# Legacy Procurement Practices Shape How U.S. Cities Govern AI: Understanding Government Employees’ Practices, Challenges, and Needs

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greeceisbn: 978-1-4503-XXXX-X/18/06copyright: noneccs: Human-centered computing Empirical studies in HCIccs: Social and professional topics Computing / technology policy
@@ 1. Introduction
Motivated by promises of increased efficiency, governments are increasingly adopting AI to automate bureaucratic workflows and assist decision-making processes that impact residents (Reisman et al., 2018; Kim et al., 2024; Chouldechova et al., 2018; Eubanks, 2018; Whitney et al., 2021; Kawakami et al., 2024a; Levy et al., 2021). Often, such public sector AI applications are not developed in house, but are purchased from external vendors through a process called public procurement (Sloane et al., 2021; Prier and McCue, 2009; Lloyd and McCue, 2004). In a 2023 statement, U.S. Senator Gary Peters shared that over half of the AI tools used by federal agencies were purchased from commercial vendors (Peters, 2023). Experts estimate that this number is even higher at lower levels of government, such as state and local governments which are less likely to have internal expertise to develop AI (Rubenstein, 2021). Thus, most public sector AI systems in use today are developed by and acquired from private vendors.
A growing number of academic and advocacy efforts have pointed out how procured AI systems have predominantly targeted narrowly defined notions of efficiency or performance, resulting in adverse effects that disproportionately impact marginalized communities (Eubanks, 2018; Roberts, 2022; Stapleton et al., 2022a; Harcourt, 2006). Facial recognition technologies, for example, have faced sustained criticism due to concerns about civil liberties, racial biases, and privacy violations (Williams, 2015; Haskins, 2021; log, 2020). Despite these concerns, in 2022, a U.S. government watchdog revealed over 20 federal contracts with private companies that either specialized in facial recognition or had awards related to its use, with a total budget exceeding \$7 million (Riley, 2022).
@@
The past few years have marked an explosion of attention directed towards imagining how AI procurement might be reformed to prepare governments to anticipate AI harms (Mulligan and Bamberger, 2019; Richardson et al., 2019; Conti-Cook, [n. d.]; Rubenstein, 2021; Hickok, 2024; gov, 2024c). However, there is a lack of empirical research to understand how AI procurement actually occurs in practice. Past scholarship has established that AI must be acquired using legacy procurement practices that apply widely to all goods and services, including pencils, school buses, and AI  (Reisman et al., 2018; Mulligan and Bamberger, 2019; Sloane et al., 2021; Rubenstein, 2021). Yet, little is known about what exactly these practices entail, how they vary across localities, and whether governments have made any changes to their practices when assessing AI specifically. Furthermore, important information about governments’ AI procurement practices is not always publicly available  (Richardson et al., 2019; Halonen, 2016).
@@
  • In Section 4, we offer an empirical account of how AI procurement occurs on-the-ground. We illustrate how legacy purchasing practices that differ across cities shape how AI is governed by establishing important infrastructure (like solicitations and contracts), determining who within a city has power over different types of AI acquisitions, and influencing the socio-technical values by which AI is assessed. By detailing the organizational factors that influence how procurement occurs, we show how cities’ purchasing practices both enable and preclude cities’ ability to address AI harms: for instance, historic norms such as cost thresholds  (Duchicela et al., 2023) allow employees to acquire low- or no-cost AI without going through the accountability measures typically associated with government purchasing. We also highlight emerging steps some cities have taken to incorporate AI-specific criteria and ethical considerations into their AI procurement processes.
  • In Section 5, we outline three key challenges that still remain to be addressed by cities’ existing (and emerging) practices. We argue that more work is urgently needed to (1) address information asymmetries between governments and vendors, (2) support cities in asking more of AI vendors, and (3) support cities in sharing and assuming ongoing responsibilities of AI governance. We discuss implications and opportunities for the FAccT community to support public sector workers in anticipating and mitigating harms caused by procured AI.
@@ 2.1. AI in the Public Sector
Governments’ increased adoption of data-driven technologies has transformed work in the public sector (Alkhatib and Bernstein, 2019; Levy et al., 2021; Engstrom and Ho, 2021; Stapleton et al., 2022b; Madan and Ashok, 2023; Kim et al., 2024). While many technologies marketed as “artificial intelligence” today bear resemblance to historical data mining technologies of the 1990s (e.g., to detect fraud (Levy et al., 2021; Engstrom and Ho, 2021)), recent years have marked a rapid increase in the number and popularity of AI tools that employ machine learning (Mulligan and Bamberger, 2019; Levy et al., 2021; Engstrom and Ho, 2021; Rubenstein, 2021). Governments’ use cases for AI, as taxonomized by (Engstrom and Ho, 2021), range from law enforcement and adjudication, such as supporting criminal sentencing (Forrest, 2021; Pruss, 2023), to applications that support government service delivery, such as chatbots to facilitate communication with residents  (Quinlan, 2024; Lecher, 2024). Like private sector workers, government employees are increasingly adopting general-purpose productivity tools like ChatGPT (David, 2024; Childress et al., 2024; Network, 2024) to do the “street-level” work of governance (Alkhatib and Bernstein, 2019), e.g., to adjudicate child welfare cases (Parmiggiani et al., 2024; Taylor, 2024), thus further blurring the lines between AI’s articulated scope and ultimate application.
@@
Scholars and activists have documented the harms resulting from public sector AI systems, with the recognition that historically oppressed communities are disproportionately subject to AI-assisted punishment (Eubanks, 2018; Stapleton et al., 2022a; Liban, 2024; Pruss et al., 2024). Public sector algorithms have faced criticism for algorithmic bias (e.g., in the case of predictive policing (Bhuiyan, 2021; in Pittsburgh, 2020)), validity (e.g., in the case of child welfare risk assessments (Coston et al., 2023; Gerchick et al., 2023)), and functionality (e.g., in the case of facial identification mismatches  (Hill, 2023)). To this end, a growing body of empirical scholarship has studied government employees’ practices, desires, and concerns surrounding governments’ use of AI systems (Veale et al., 2018; Brown et al., 2019; Saxena and Guha, 2020; Stapleton et al., 2022a; Kuo et al., 2023; Cheng et al., 2022; Kawakami et al., 2022; Ziosi and Pruss, 2024). Much research has focused on the perspectives of frontline workers (e.g., social workers (Kawakami et al., 2022)) who consume AI outputs. Studies have characterized how AI systems’ “misalignment” with policy objectives and workers’ ethical values can result in harms to marginalized communities (Kawakami et al., 2022; Stapleton et al., 2022a; Robertson et al., 2021), highlighting implications for technology design. Other past work illuminates the organizational factors that facilitate critical decisions about AI adoption and governance (Veale et al., 2018; Kawakami et al., 2022; Pruss, 2023; Kawakami et al., 2024a, b), realizing that governments “should not be treated as monolithic entities” (Kawakami et al., 2024a), and that workers at varying levels of institutional power often hold varying perspectives about the appropriateness and ethical implications of AI tools.
@@ 2.2. Governing AI through Public Procurement
The term “public procurement” generally refers to the processes governments use to bring in goods and services that are developed externally (Sloane et al., 2021; Prier and McCue, 2009; Lloyd and McCue, 2004), often involving paid transactions with third-party organizations. [^1] We note that as described by past work (Prier and McCue, 2009), there is no single precise agreed-upon definition for what it meant by the term “public procurement” – rather, the definition is “muddled” and varies across contexts. See Appendix A for a more detailed discussion of definitions. In this work, our core focus is on studying public procurement practices in the United States, particularly for local (city) governments. Public administration scholars have characterized how procurement laws, organizational structures, and activities vary between levels of government and localities (Coggburn, 2003; Lloyd and McCue, 2004; Duchicela et al., 2023).
@@
A growing coalition of policymakers, scholars, and civil society groups have called on governments to implement AI-specific procurement guidelines (uk_, 2020; Richardson et al., 2019; Conti-Cook, [n. d.]; Rubenstein, 2021; Hickok, 2024). Governments must be prepared, scholars argue (Mulligan and Bamberger, 2019; Rubenstein, 2021; de la Vallee et al., 2024), to interrogate the value-laden trade-offs embedded by design decisions such as the vendor’s choice of training dataset, optimization function, and predictive target. In response, several groups have explored how to embed these and other ethical considerations within cities’ existing procurement processes  (gov, 2024c; Richardson et al., 2019; Rubenstein, 2021; Sharma and Adler, 2024). These efforts have produced practical guidance and readily adoptable resources for government employees, such as guidelines, tools, vendor repositories, and templates to guide AI procurement practices (e.g., (Richardson et al., 2019; wef, 2020; Conti-Cook, [n. d.]; Dotan et al., 2023; Rubenstein, 2021; Brauneis and Goodman, 2018; Project, 2024)). However, there is very little work that examines whether and how these tools help government employees in practice.
@@
Many AI procurement reform initiatives draw from a long history of governments adapting their purchasing practices to enact social change  (Rubenstein, 2021; Grandia and Meehan, 2017; McCrudden, 2007), e.g., by creating processes that prioritize minority-owned businesses (McCrudden, 2007), enable public oversight over government surveillance (Crump, 2016; Young et al., 2019), or incentivize other ethical behaviors, such as sustainability (Lăzăroiu et al., 2020; Varga, 2021). These initiatives call attention to the purchasing power governments hold “not as regulators, but as participants” in a free market  (McCrudden, 2007). Drawing from this tradition, scholars point to the “gatekeeping role” of public procurement in deciding which AI systems are (and are not) purchased (Sloane et al., 2021; Oluka et al., 2022; Conti-Cook, [n. d.]; Hickok, 2024). The ultimate hope expressed is that governments’ interactions with vendors can “change markets” (wef, 2020): that governments can exercise their purchasing power to advocate for AI vendors to prioritize social values such as fairness, accountability, and transparency  (Sharma and Adler, 2024). Given this potential role of procurement in AI governance, there have been a small number of research efforts that have uncovered challenges faced by AI procurement practitioners in Kenya and Uganda  (Oluka et al., 2022) and the US federal government  (Autio et al., 2023), such as knowledge gaps in public sector workers’ understanding of AI risks or an absence of “clearly defined standards” for AI governance. However, what remains less known is the current practices that governments follow when acquiring AI, e.g., the assessment criteria governments apply, how purchasing decisions are made, and who holds the power to make them.
@@
This paper contributes to these ongoing discussions by illuminating the organizational realities in which AI procurement occurs. We provide an empirical account of the emerging actions that U.S. local governments have taken to revise their AI procurement processes on the ground. Given this context, in Section 5 we revisit several critical challenges surfaced by past scholarship, such as knowledge gaps in government employees’ preparedness to anticipate AI harms (Zick et al., 2024; Oluka et al., 2022; Engstrom and Ho, 2021), governments’ lack of leverage with AI vendors (Kawakami et al., 2024a; Hickok, 2024), and uncertainty about how to share and delegate responsibilities of AI governance (aim, 2024b; Groves et al., 2024). We ground our analysis using examples of real-world procurements to reflect on both the progress cities have made and obstacles that remain to be overcome.
@@ 3. Methodology
To understand the organizational processes that shape how cities govern AI, we focus on the perspectives of municipal employees. We conducted semi-structured interviews over six months from December 2023 to June 2024. Our research coincided with a landmark year of government action on AI procurement, including the release and adoption of the U.S. “AI M-Memo” (aim, 2024a; Shivaram, 2024), which instituted governance requirements for federal agencies’ use of AI [^2] The Biden AI M-Memo has since been revoked by the new administration  (O’Brien, 2025). , and the formation of the Government AI Coalition (gov, 2024d; Edinger, 2023; Rispens, 2024), a grassroots coalition founded to “give local governments a voice in shaping the future of AI”. As a result, our study took place at a moment when governments across the U.S. were just beginning to stand up and implement new governance processes for procured AI.
Recruitment. Our recruitment process aimed to capture a diverse set of employees’ perspectives on the AI procurement process, across multiple cities. Many cities required employees to complete a formal approval process to participate in research, so we organized our recruitment efforts by city. Once we received approval, we used snowball sampling to ask our initial contacts within a city to introduce us to other employees whose present role was involved with technology procurement or governance. As shown in Table 1, most participants worked in technology-focused roles in their city’s IT or Innovation departments. We also spoke with specialists in vendor relations and procurement, and one HR specialist who had conducted organizational training on AI. Participants included both leaders who made decisions on behalf of their department, and workers whose day-to-day responsibilities involved managing AI procurement.
We began the study by reaching out to contacts in our professional networks in four U.S. cities, who introduced us to contacts in three additional cities. From the contacts we were given, we intentionally invited cities that represented a wide range of regions and maturity surrounding AI (e.g., whether or not they had adopted any public-facing AI policies). Although this approach has limitations, we learned that establishing trust through shared connections was important for government employees to feel comfortable speaking openly with academic researchers. For example, one participant shared that their department was hesitant to participate in research studies led by someone they were unfamiliar with. We invited eight total cities to participate, and seven agreed. Participating cities represented all four major regions (Northeast, West, Midwest, and South) defined by the U.S. Census Bureau (cen, 2024).
@@
P15 Ryan Information Technology Large Chief Technology Officer
P16 Adrian Innovation Large IT Policy Director
@@
Table 1. An anonymous description of participating municipal employees. Titles were modified to preserve anonymity. Small cities have under 200,000 residents, medium cities have 200,000 - 500,000 residents, and large cities have over 500,000 residents.
@@
Semi-structured interviews. Following Veale et al. (2018), we adopted a semi-structured interview approach to allow for flexibility in discussions. This allowed participants to spend more time discussing the phases of the procurement process that were closest to their role and expertise. The interviews ranged from 60 to 90 minutes and our protocol included three sections. First, we asked participants about their current role and work responsibilities relating to AI. We defined “AI” for participants using the OECD definition of “any machine-based system that can make predictions, recommendations, or decisions” (for Economic Cooperation and , OECD), and provided examples of qualifying systems. Second, we asked participants to walk through how an example procurement involving AI would occur in their city, paying particular attention to any differences between a standard technology procurement. Third, we asked participants to reflect more deeply on their perceived challenges, needs, and desires to improve the AI procurement process. We expressed to participants that we were interested in understanding potential risks or societal impacts of the AI systems they procured, but did not steer participants towards discussing particular types of impacts, as we aimed to understand how participants conceptualized risk and harm in their own words. The study was approved by a university IRB. We include our complete interview protocol in Appendix D and discuss ethics and participant safety considerations in Section 7.1.
@@
Data analysis. We collected 23 hours of interview audio which were transcribed and coded by four team members. We adopted a bottom-up thematic analysis approach (Braun and Clarke, 2006) to analyze interview transcripts. Each transcript was first open-coded by two authors, who met to discuss each transcript and resolve any differences in interpretation (McDonald et al., 2019). In total, we created 305 unique codes, such as “proprietariness” or “indemnification from AI harms”. We then grouped these codes into themes corresponding to cities’ purchasing practices, such as “alternative procurement pathways” and “AI risk assessments”, and challenges faced by employees such as “lack of leverage with AI vendors”. We also complemented our interview findings by conducting a document analysis  (Karppinen and Moe, 2012) of cities’ written AI policies or procurement materials, such as solicitation or contracting templates that were referenced in interviews. When selecting our final set of themes, we aimed to emphasize findings that contributed new insights to the existing literature on AI and procurement.

### v3

@@ Abstract.
journalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732049isbn: 979-8-4007-1482-5/2025/06copyright: noneccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Motivated by promises of increased efficiency, governments worldwide are increasingly adopting AI to automate bureaucratic workflows and assist decision-making processes that impact residents (Reisman et al., 2018; Kim et al., 2024; Chouldechova et al., 2018; Eubanks, 2018; Whitney et al., 2021; Kawakami et al., 2024a; Levy et al., 2021). We situate our study in the United States, where such public sector AI applications are often not developed in house, but are instead purchased from external vendors through a process called public procurement (Sloane et al., 2021; Prier and McCue, 2009; Lloyd and McCue, 2004). In a 2023 statement, U.S. Senator Gary Peters shared that over half of the AI tools used by federal agencies were purchased from commercial vendors (Peters, 2023). Scholars estimate that this number is even higher at lower levels of government, such as U.S. state and local governments which are less likely to have internal expertise to develop AI (Rubenstein, 2021). Thus, most AI systems used by U.S. governments today are developed and acquired from private vendors.
A growing number of academic and advocacy efforts have pointed out how procured AI systems have predominantly targeted narrowly defined notions of efficiency or performance, resulting in adverse effects that disproportionately impact marginalized communities (Eubanks, 2018; Roberts, 2022; Stapleton et al., 2022a; Harcourt, 2006). Facial recognition technologies, for example, have faced sustained criticism due to concerns about civil liberties, racial biases, and privacy violations (Williams, 2015; Haskins, 2021; Logic Magazine(2020), s). Despite these concerns, in 2022, a U.S. government watchdog revealed over 20 federal contracts with private companies that either specialized in facial recognition or had awards related to its use, with a total budget exceeding \$7 million (Riley, 2022).
@@
The past few years have marked an explosion of attention directed towards imagining how AI procurement might be reformed to prepare governments to anticipate AI harms (Mulligan and Bamberger, 2019; Richardson et al., 2019; Conti-Cook, [n. d.]; Rubenstein, 2021; Hickok, 2024; The Government AI Coalition, 2024c). However, there is a lack of empirical research to understand how AI procurement actually occurs in practice. Past scholarship has established that governments must acquire AI using legacy procurement practices that apply widely to all goods and services, including pencils, school buses, and AI  (Reisman et al., 2018; Mulligan and Bamberger, 2019; Sloane et al., 2021; Rubenstein, 2021). Yet, little is known about what exactly these practices entail, how they vary across localities, and whether governments have made any changes to their practices when assessing AI specifically. Furthermore, important information about governments’ AI procurement practices is not always publicly available  (Richardson et al., 2019; Halonen, 2016).
@@
In Section 4, we offer an empirical account of how AI procurement occurs on-the-ground. By detailing the organizational factors that influence how procurement occurs, we show how cities’ purchasing practices both enable and preclude cities’ ability to address AI harms: for instance, historic norms such as cost thresholds  (Duchicela et al., 2023) allow employees to acquire low- or no-cost AI without going through the accountability measures typically associated with government purchasing. We also highlight emerging steps some cities have taken to incorporate AI-specific criteria and ethical considerations into their AI procurement processes.
In Section 5, we outline three key challenges that still remain to be addressed by cities’ existing (and emerging) practices. We argue that more work is urgently needed to (1) address information asymmetries between governments and vendors, (2) support cities in asking more of AI vendors, and (3) support cities in sharing and assuming ongoing responsibilities of AI governance. We discuss implications and opportunities for the FAccT community to support public sector workers in mitigating harms caused by procured AI.
Our research team has developed an accompanying white paper to this manuscript that translates our research findings into actionable recommendations for local government employees  (Silva et al., 2025). The white paper is available online (link).
@@ 2.1. AI in the Public Sector
Governments’ increased adoption of data-driven technologies has transformed work in the public sector (Alkhatib and Bernstein, 2019; Levy et al., 2021; Engstrom and Ho, 2021; Stapleton et al., 2022b; Madan and Ashok, 2023; Kim et al., 2024). While many technologies marketed as “artificial intelligence” today bear resemblance to historical data mining technologies of the 1990s (e.g., to detect fraud (Levy et al., 2021; Engstrom and Ho, 2021)), recent years have marked a rapid increase in the number and popularity of AI tools that employ machine learning (Mulligan and Bamberger, 2019; Levy et al., 2021; Engstrom and Ho, 2021; Rubenstein, 2021). Governments’ use cases for AI, as taxonomized by (Engstrom and Ho, 2021), range from law enforcement and adjudication, such as supporting criminal sentencing (Forrest, 2021; Pruss, 2023), to applications that support government service delivery, such as chatbots to facilitate communication with residents  (Quinlan, 2024; Lecher, 2024). Like private sector workers, government employees are increasingly adopting general-purpose productivity tools such as ChatGPT (David, 2024; Childress et al., 2024; MetroLab Network, 2024) to do the “street-level” work of governance (Alkhatib and Bernstein, 2019), e.g., to adjudicate child welfare cases (Parmiggiani et al., 2024; Taylor, 2024), thus further blurring the lines between AI’s articulated scope and ultimate application.
@@
Scholars and activists have documented the harms resulting from public sector AI systems, with the recognition that historically oppressed communities are disproportionately subject to AI-assisted punishment (Eubanks, 2018; Stapleton et al., 2022a; Liban, 2024; Pruss et al., 2024). Public sector algorithms have faced criticism for algorithmic bias (e.g., in the case of predictive policing (Bhuiyan, 2021; The Coalition Against Predictive Policing in Pittsburgh, 2020)), lack of validity (e.g., in the case of child welfare risk assessments (Coston et al., 2023; Gerchick et al., 2023)), and lack of functionality (e.g., in the case of facial identification mismatches  (Hill, 2023)). In response, a growing body of empirical scholarship has studied government employees’ practices, desires, and concerns surrounding governments’ use of AI systems (Veale et al., 2018; Brown et al., 2019; Saxena and Guha, 2020; Stapleton et al., 2022a; Kuo et al., 2023; Cheng et al., 2022; Kawakami et al., 2022; Ziosi and Pruss, 2024). Much research has focused on the perspectives of frontline workers (e.g., social workers (Kawakami et al., 2022)) who consume AI outputs. Studies have characterized how AI systems’ “misalignment” with policy objectives and workers’ ethical values can result in harms to marginalized communities (Krafft et al., 2021; Kawakami et al., 2022; Stapleton et al., 2022a; Robertson et al., 2021), highlighting implications for technology design. Other past work illuminates the organizational factors that facilitate critical decisions about AI adoption and governance (Krafft et al., 2021; Veale et al., 2018; Kawakami et al., 2022; Pruss, 2023; Kawakami et al., 2024a, c), recognizing that governments “should not be treated as monolithic entities” (Kawakami et al., 2024a), and that workers at varying levels of institutional power often hold varying perspectives about the appropriateness and ethical implications of AI tools.
@@ 2.2. Governing AI through Public Procurement
The term “public procurement” generally refers to the processes governments use to bring in goods and services that are developed externally (Sloane et al., 2021; Prier and McCue, 2009; Lloyd and McCue, 2004), often involving paid transactions with third-party organizations. [^1] We note that as described by past work (Prier and McCue, 2009), there is no single precise agreed-upon definition for what it meant by the term “public procurement” – rather, the definition is “muddled” and varies across contexts. See Appendix A for a more detailed discussion of definitions. In this work, our core focus is on studying public procurement practices in the United States, particularly for municipal (city) governments. Public administration scholars have characterized how procurement laws, organizational structures, and activities generally vary between levels of government and localities (Coggburn, 2003; Lloyd and McCue, 2004; Duchicela et al., 2023).
@@
A growing coalition of policymakers, scholars, and civil society groups have called on governments to implement AI-specific procurement guidelines (UK Government, 2020; Richardson et al., 2019; Conti-Cook, [n. d.]; Rubenstein, 2021; Hickok, 2024). Governments must be prepared, scholars argue (Mulligan and Bamberger, 2019; Rubenstein, 2021; de la Vallee et al., 2024), to interrogate the value-laden trade-offs embedded by design decisions such as the vendor’s choice of training dataset, optimization function, and predictive target. In response, several groups have explored how to embed these and other ethical considerations within cities’ existing procurement processes  (The Government AI Coalition, 2024c; Richardson et al., 2019; Rubenstein, 2021; Sharma and Adler, 2024). These efforts have produced practical guidance and readily adoptable resources for government employees, such as guidelines, tools, vendor repositories, and templates to guide AI procurement practices (e.g., (Richardson et al., 2019; World Economic Forum, 2020; Conti-Cook, [n. d.]; Dotan et al., 2023; Rubenstein, 2021; Brauneis and Goodman, 2018; EdTech Equity Project, 2024)). However, there is very little work that examines whether and how these tools help government employees in practice.
@@
Many AI procurement reform initiatives draw from a long history of governments adapting their purchasing practices to enact social change  (Rubenstein, 2021; Grandia and Meehan, 2017; McCrudden, 2007), e.g., by creating processes that prioritize minority-owned businesses (McCrudden, 2007), enable public oversight over government surveillance (Crump, 2016; Young et al., 2019), or incentivize other ethical behaviors, such as sustainability (Lăzăroiu et al., 2020; Varga, 2021). These initiatives call attention to the purchasing power governments hold “not as regulators, but as participants” in a free market  (McCrudden, 2007). Drawing from this tradition, scholars point to the “gatekeeping role” of public procurement in deciding which AI systems are (and are not) purchased (Sloane et al., 2021; Oluka et al., 2022; Conti-Cook, [n. d.]; Hickok, 2024). The ultimate hope expressed is that governments’ interactions with vendors can “change markets” (World Economic Forum, 2020): that governments can exercise their purchasing power to advocate for AI vendors to prioritize social values such as fairness, accountability, and transparency  (Sharma and Adler, 2024).
Given this potential role of procurement in AI governance, there have been a small number of research efforts that have uncovered challenges faced by AI procurement practitioners in Kenya and Uganda  (Oluka et al., 2022) and the US federal government  (Autio et al., 2023), such as knowledge gaps in workers’ understanding of AI risks or an absence of “clearly defined standards” for AI governance. However, what remains less known is the current practices that governments follow when acquiring AI, e.g., the assessment criteria governments apply, how purchasing decisions are made, and who holds the power to make them.
@@
This paper contributes to these ongoing discussions by illuminating the organizational realities in which AI procurement occurs. We provide an empirical account of the emerging actions that U.S. local governments have taken to revise their AI procurement processes on the ground. Given this context, in Section 5 we revisit several critical challenges surfaced by past scholarship, such as knowledge gaps in government employees’ preparedness to anticipate AI harms (Zick et al., 2024; Oluka et al., 2022; Engstrom and Ho, 2021), governments’ lack of leverage with AI vendors (Kawakami et al., 2024a; Hickok, 2024), and uncertainty about how to share and delegate responsibilities of AI governance (US Office of Management and Budget, 2024; Groves et al., 2024). We ground our analysis using examples of real-world procurements to reflect on both the progress cities have made and obstacles that remain to be overcome.
@@ 3. Methodology
To understand the organizational processes that shape how cities govern AI, we focus on the perspectives of municipal employees. We conducted semi-structured interviews over six months from December 2023 to June 2024. Our research coincided with a landmark year of government action on AI procurement, including the release and adoption of the U.S. “AI M-Memo” (Executive Office of the President, 2024; Shivaram, 2024), which instituted governance requirements for federal agencies’ use of AI, [^2] The Biden AI M-Memo has since been revoked by the Trump administration  (O’Brien, 2025). and the formation of the Government AI Coalition (Edinger, 2023; Rispens, 2024), a grassroots coalition founded to “give local governments a voice in shaping the future of AI.” As a result, our study took place at a moment when governments across the U.S. were just beginning to stand up and implement new governance processes for procured AI.
Recruitment. Our recruitment process aimed to capture a diverse set of employees’ perspectives on the AI procurement process, across multiple cities. Many cities required employees to receive approval to participate in research, so we organized our recruitment efforts by city. Once we received approval, we used snowball sampling to ask our initial contacts within a city to introduce us to other employees whose present role was involved with technology procurement or governance. As shown in Table 1, most participants worked in technology-focused roles in their city’s IT or Innovation departments. We also spoke with specialists in vendor relations and procurement, and one HR specialist who had conducted organizational training on AI. Participants included both leaders who made decisions on behalf of their department, and workers whose day-to-day responsibilities involved managing AI procurement.
We began the study by reaching out to contacts in our professional networks in four U.S. cities, who introduced us to contacts in three additional cities. From the contacts we were given, we intentionally invited cities that represented a wide range of regions and maturity surrounding AI (e.g., whether or not they had adopted any public-facing AI policies). Although this approach has limitations, we learned that establishing trust through shared connections was important for government employees to feel comfortable speaking openly with academic researchers. For example, one participant shared that their department was hesitant to participate in research studies led by someone they were unfamiliar with. We invited eight total cities to participate, and seven agreed. Participating cities represented all four major regions (Northeast, West, Midwest, and South) defined by the U.S. Census Bureau (U.S. Census Bureau, 2024).
@@
P15 Buck Information Technology Large Chief Technology Officer
P16 Adrian Innovation Large IT Director
@@
Table 1. An anonymous description of participating municipal employees. Titles were modified to preserve anonymity. Small cities have under 200,000 residents, medium cities have 200,000 - 500,000 residents, and large cities have over 500,000 residents. Participants were invited to choose their own pseudonym.
Semi-structured interviews. Following Veale et al. (2018), we adopted a semi-structured interview approach to allow for flexibility in discussions. This allowed participants to spend more time discussing the phases of the procurement process that were closest to their role and expertise. The interviews ranged from 60 to 90 minutes and our protocol included three sections.
@@
First, we asked participants about their current role and work responsibilities relating to AI. We defined “AI” for participants using the OECD definition of “any machine-based system that can make predictions, recommendations, or decisions” (The Organization for Economic Cooperation and Development (2019), OECD), and provided examples of qualifying systems. We chose the OECD definition to be intentionally inclusive of a wide set of predictive technologies that pose risks to residents’ rights and safety, following a best practice from civil society  (de la Vallee et al., 2024). Second, we asked participants to walk through how an example procurement involving AI would occur in their city, paying particular attention to any differences between a standard technology procurement. Third, we asked participants to reflect more deeply on their perceived challenges, needs, and desires to improve the AI procurement process. We expressed to participants that we were interested in understanding potential risks or societal impacts of the AI systems they procured, but did not steer participants towards discussing particular types of impacts, as we aimed to understand how participants conceptualized risk and harm in their own words. The study was approved by a university IRB. We include our complete interview protocol in Appendix E and discuss ethics and participant safety considerations in Section 7.1.
@@
Data analysis. We collected 23 hours of interview audio which were transcribed and coded by four team members. We adopted a bottom-up thematic analysis approach (Braun and Clarke, 2006) to analyze interview transcripts. Each transcript was first open-coded by two authors, who met to discuss each transcript and resolve any differences in interpretation (McDonald et al., 2019). In total, we created 305 unique codes, such as “proprietariness” or “indemnification from AI harms”. We then grouped these codes into themes corresponding to cities’ purchasing practices, such as “alternative procurement pathways”, and challenges faced by employees, such as “lack of leverage with AI vendors”. We complemented interview findings by conducting a document analysis  (Karppinen and Moe, 2012) of cities’ written AI policies or procurement materials. When available, we collected and read cities’ procurement handbooks, AI policies, AI risk assessment instruments, solicitation templates, and contracting templates. [^3] Such documents are not always publicly available, and were sometimes provided by request by our interviewees. We primarily used supporting documents to corroborate findings and themes observed from interviews: for example, after learning from interviews that cost thresholds played a significant role in AI governance, we consulted each city’s procurement handbook to ensure we understood the cost thresholds applicable for each city. When selecting our final set of themes, we aimed to emphasize findings that contributed new insights to the existing literature on AI and procurement.